A normal carpenter's ruler marked off in sixteenths. What would be the values at each of the marks in decimal form?
Answer by Zaraisty(1)

1/16 = .0625
1/8 = .125
3/16 = .1875
1/4 = .25
5/16 = .3125
3/8 = .3750
7/16 = .4375
1/2 = .5
9/16 = .5625
5/8 = .625
11/16 = .6875
3/4 = .75
13/16 = .8125
7/8 = .875
15/16 = .9375
1 =1